OpenStudy (anonymous):

in africa there is a high population of cicada (a type of insect) every 17 years. in the same part of Africa there is a high population of lizards every 12 years.There was a high population of both cicadas and lizards in year 2001. In what year will there next be a large population of both cicadas and lizards ?

OpenStudy (anonymous):

that wld be lcd of 12 and 17

OpenStudy (anonymous):

there is another type of cicada that appears every 13 years

OpenStudy (anonymous):

so they only appear together every 13* 17 tears

OpenStudy (anonymous):

so whats the next year after 2001 that there will be a large population of both ?

OpenStudy (anonymous):

righ tadd on lcd of 12 and 17 12*17 = 204 try 102 - no lcd = 204 so 2001 + 204 = 2205

OpenStudy (anonymous):

year 2205